Circular No. 6611

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GS 1826–238 P. Ubertini, A. Bazzano, M. Cocchi, and L. Natalucci, Istituto di As-

trofisica Spaziale, CNR, Frascati; J. Heise, R. Jager, and J. in ’t Zand, Space Research Organisation of the Netherlands (SRON), Utrecht; J. M. Muller and M. Smith, SRON and BeppoSAX Scientific Data Center, Rome; and G. Celidonio, A. Coletta, R. Ricci, P. Giommi, D. Ricci, M. Capalbi, M. T. Menna, and S. Rebecchi, BeppoSAX, Rome, report: “The Ginga transient GS 1826–24 was detected beginning on Mar. 29.61 UT at a flux on the order of 12 mCrab in the range 1.8–8.3 keV with the Wide Field Camera onboard the BeppoSax satellite, during a deep galactic-center observation. The preliminary source position is α = 18^{h}29^{m}28^{s}, δ = −23^{o}48′.0 (equinox 2000.0), with a 3′ error radius. During the follow-up monitoring of the source, at least three bursts were detected, on Mar. 30.092 (peak flux 350 mCrab), 31.694 (520 mCrab), and 31.943 (480 mCrab). All bursts show a duration time of ∼ 150 s. On the basis of the bursting behavior, we suggest that the source is a neutron-star system (as possibly first indicated by the OSSE high-energy spectrum; Strickman et al. 1996, A.Ap. Suppl. Ser. 120, 217).”

GRB 970228 S. Klose and B. Stecklum, Thüringer Landessternwarte Tautenburg,

and R. Tuffs, Max-Planck-Institut für Kernphysik, Heidelberg, write: “The error box of GRB 970228 was surveyed with the Calar Alto 3.5-m telescope (f/10) on Mar. 17.8 UT using the near-infrared camera MAGIC (0′′.81 pixels, seeing < 1′′.5, photometric sky). At the position of the optical transient reported on IAUC 6584, a source is detected (5σ) in the J band (magnitudes J = 21.0, H > 20.0, K > 19.5). We cannot decide, however, whether this is the reported underlying galaxy (Groot et al. and Metzger et al., IAUC 6588) or the fading infrared counterpart of the optical transient (or both). The J-band image can be accessed via http://www.tls-

tautenburg.de/research.html.”

COMET C/1997 D1 (MUELLER) Visual m_{1} estimates: Mar. 1.80 UT, 14.0 (K. Hornoch, Lelekovice,

Czech Republic, 0.35-m reflector); 4.27, 13.4 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 28.91, 13.6 (J. Carvajal, Teruel, Spain, 0.44-m reflector); Apr. 2.92, 13.4 (M. Meyer, Frauenstein, Germany, 0.25-m reflector).

1997 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green